# Autonomous Driving in Unstructured Environments: How Far Have We Come?

### v3

@@ II-B3 Place Recognition / Relocalization
Tian et al. [113] Scan Context Multi-scene LiDAR
Ou et al. [114] Spatial Binary Pattern Orchard LiDAR
Yuan et al. [115] Binary and Triangle Combined Descriptor Multi-scene LiDAR
Ramezani et al. [116] Global and Local Descriptor Natural Scene LiDAR/Camera
Oh et al. [117] Diverse Descriptors for Evaluation Dense Forest IMU/LiDAR
Barros et al. [118] PointNetPGAP-based Descriptor Horticulture LiDAR
@@
Place Recognition (PR) technology enables vehicles to identify previously visited environments, playing a crucial role in autonomous driving. This technology can be utilized to reduce drift errors of SLAM systems, ensuring reliable localization and globally consistent map construction. In unstructured environments, the PR task is particularly challenging due to the dynamic nature and sparse, diverse features of the surroundings. Such conditions often render methods [119, 120, 121] designed for urban environments ineffective. Currently, there are some studies that focus on PR technology in unstructured environments, as shown in Table IV.
@@
Tian et al. [113] propose STV-SC, which enhances the scan context [119] algorithm by incorporating segmentation and temporal verification. This method includes a range image-based 3D point segmentation algorithm to extract key structured information and a three-stage process for loop detection, significantly improving recall rates compared to state-of-the-art algorithms. Ou et al. [114] introduce a novel LiDAR-based PR algorithm that utilizes the spatial binary pattern to encode 3D spatial information into an eight-bit binary pattern. The two-stage hierarchical re-identification process with attention score maps further enhances the method’s effectiveness in orchard and public datasets. Another innovative method is [115], which combines binary and triangle descriptors to achieve robust 3D place recognition. This method forms triangles from keypoints and uses their side lengths to create a global descriptor invariant to pose changes, significantly improving precision and adaptability, especially in scenarios with large viewpoint variations.
@@
Further advancements in this domain involve integrating deep learning techniques for improved robustness in challenging scenarios. Ramezani et al. [116] present a method that integrates LiDAR and image data through a self-supervised 2D-3D feature matching process, leveraging deep learning for LiDAR feature extraction and relative pose estimation between point clouds. This approach improves robustness and accuracy in multi-robot relocalization tasks in unstructured environments. Oh et al. [117] evaluate handcrafted and learning-based models in dense forest environments, integrating the Logg3dNet model into a robust 6-DoF pose estimation system.
@@
Barros et al. [118] introduce PointNetPGAP-SLC, a model combining a global average pooling aggregator and a pairwise feature interaction aggregator to enhance descriptor generation. The Segment-Level Consistency (SLC) model used during training improves the robustness of descriptors. Evaluations on the dataset demonstrate that PointNetPGAP improves PR performance in horticultural settings by addressing descriptor ambiguity and leveraging the SLC model for enhanced training.
@@
Moreover, the development of specific datasets and evaluation frameworks has been crucial in advancing PR techniques in natural environments. Hausler et al. [122] address the lack of benchmarks for natural, unstructured environments by introducing the Wild-Places [63] and WildScenes [59] datasets. These datasets enable comprehensive evaluation and development of localization and mapping algorithms in challenging unstructured settings, bridging the gap between urban and natural environments.
II-C Odometry-based Methods
@@
Marks et al. [12] present Gamma-SLAM, an online stereo visual SLAM algorithm that uses a Rao-Blackwellized particle filter to estimate joint posterior distributions over poses and maps. This method incorporates visual odometry for accurate proposal distributions and represents maps through a Cartesian grid, maintaining a posterior distribution over elevation variance in each cell to effectively capture terrain features in unstructured environments. Zhang et al. [13] introduce LOAM, a real-time LiDAR odometry and mapping method that divides the problem into high-frequency odometry and low-frequency mapping algorithms. This approach ensures low drift and low computational complexity, making it suitable for real-time applications. Extending this concept, Zhang and Singh [101] combine visual and LiDAR odometry (V-LOAM) to enhance localization and mapping by integrating high-frequency visual odometry for initial motion estimation and low-frequency LiDAR odometry for refinement, achieving robust performance even under aggressive motion and poor visual conditions.
@@
Further advancements in this domain include methods that focus on optimizing computational efficiency and robustness in varying terrains. Shan and Englot [97] introduce LeGO-LOAM, a lightweight and ground-optimized LiDAR odometry and mapping method that achieves real-time six degree-of-freedom pose estimation using low-power embedded systems. This method leverages the ground plane for segmentation and optimization, enhancing accuracy and computational efficiency. Based on [97, 102], Si et al. [103] establish fixed and generative road surface models to constrain the vertical drift errors of autonomous vehicles in outdoor scenarios. Since ground point clouds can be continuously and stably obtained, this ground-constrained approach could also be suitable for feature-sparse off-road environments.
@@
Giubilato et al. [104] present GPGM-SLAM, which leverages Gaussian process gradient maps to improve place recognition robustness and facilitate loop closure detection in planetary analog environments, demonstrating superior performance compared to traditional visual SLAM methods. Additionally, Chen et al. [15] propose a lightweight and efficient LiDAR odometry method that handles dense point clouds without significant preprocessing, using an adaptive keyframing system and a custom iterative closest point solver to reduce computational overhead.
@@
Yuan et al. [14] introduce a novel voxel mapping method designed for LiDAR odometry, incorporating uncertainties from LiDAR measurement noise and pose estimation errors, and demonstrating superior accuracy and efficiency in various environments. Wang et al. [105] present a LiDAR-based SLAM algorithm that leverages neighborhood information constraints to enhance feature point discrimination and prevent mis-registration in rugged terrains, showing significant improvements in localization and mapping accuracy. Tuna et al. [106] introduce X-ICP, which incorporates localizability-aware constraints in LiDAR registration to ensure reliable pose updates in geometrically uninformative environments. Dai et al. [107] propose an intensity-enhanced LiDAR SLAM framework that leverages both geometric and intensity information from LiDAR data to improve localization accuracy in sparse feature environments.
@@
The strategy of multi-sensor fusion has gained significant attention in recent years as an important means to effectively enhance the reliability and robustness of SLAM technology. For instance, Xue et al. [72] integrate IMU, wheel encoder, and LiDAR odometry to estimate vehicle motion, using IMU data to correct LiDAR scan distortions and provide better initial guesses for LiDAR odometry. This method demonstrates improved accuracy and high-frequency localization in various environments without prior information. Similarly, Zhang and Singh [108] present a modularized data processing pipeline that integrates range, vision, and inertial sensing to estimate motion and construct maps through multilayer optimization, ensuring high accuracy and low drift even under aggressive motion and environmental changes.
@@
Palieri et al. [109] present LOCUS, a multi-sensor LiDAR-centric solution that integrates multiple sensing modalities and employs a robust scan-matching algorithm to ensure continuous operation despite sensor failures. This system demonstrates superior localization accuracy and robustness, particularly in challenging settings such as the DARPA Subterranean Challenge [64]. Additionally, Shan et al. [110] introduce LVI-SAM, a tightly-coupled LiDAR-visual-inertial odometry framework that leverages a factor graph for global optimization. This method enhances visual features with LiDAR depth information and employs a robust failure detection mechanism, achieving high accuracy and robustness in GPS-denied and feature-sparse environments.
@@
Lee et al. [111] present Switch-SLAM, which uses a switching structure to transition between LiDAR and visual odometry based on detected degeneracies, ensuring robust and accurate localization across various challenging environments without the need for heuristic tuning. Additionally, Lan et al. [112] introduce a multi-sensor fusion localization system that integrates LiDAR, vision, and IMU data to achieve high accuracy and robustness in GNSS-denied environments, dynamically adjusting fusion weights in response to sensor degradation and demonstrating superior performance in urban areas, tunnels, and unstructured mountainous roads.
III Environmental Perception
@@
III-A Challenges of Environmental Perception in Unstructured Environment
@@
III-B Traversability Estimation
@@
III-B1 LiDAR-based Methods
Traversability estimation methods based on geometric information focus on extracting and analyzing geometric features from LiDAR point cloud to classify the passability of the environment. These methods can be further categorized into feature-based methods and model-based methods, as shown in Table V.
@@ III-B1 LiDAR-based Methods
Feature-based methods for analyzing traversability primarily rely on extracting characteristics from LiDAR point cloud to assess terrain passability. A notable example is the Min-Max elevation mapping technique introduced by Thrun et al. [16]. This approach begins by projecting the 3D LiDAR point cloud into a 2D grid map. The method evaluates the traversability of each grid cell by examining the height variations of points within that cell. Besides height differences, other parameters such as average height [124], covariance [124], slope [125], and roughness[126] can also be employed to analyze traversability. These techniques are straightforward to implement and have relatively low computational demands. However, they often lack robustness and can struggle in adapting to environments characterized by rough terrain.
@@
To enhance robustness in traversability analysis, various machine learning algorithms can be employed. In [127], the Expectation Maximization (EM) algorithm is utilized to fit a Gaussian Mixture Model (GMM) based on labeled statistical features, categorizing terrains in forest environments into three distinct groups. A Support Vector Machine (SVM) classifier, as described in [128], distinguishes between traversable and non-traversable regions, with each grid cell represented in an NDT traversability map (NDT-TM). In unstructured environments, a random forest classifier has been developed in [129] to identify traversable areas. Furthermore, Lee and Chung [130] detail the training of a Multi-Layer Perceptron (MLP) model for traversability mapping using a self-training algorithm. More recently, Guan et al. [131] combined geometric features, such as slope and step height, with semantic attributes to enhance traversability analysis.
@@
With the rise of deep learning, many researchers have focused on applying it for feature extraction and traversability analysis. To effectively leverage Convolutional Neural Networks (CNNs), two primary strategies have been proposed for processing unordered 3D point clouds. The first strategy involves re-encoding point clouds to create a dense input tensor suitable for CNNs. For instance, some studies have transformed point clouds into elevation maps [132], 2D Birds-Eye-View (BEV) maps [133], or range images [134], enabling CNNs to detect traversable regions in an end-to-end manner. The second strategy employs architectures like PointNet [135], designed specifically for unordered 3D point sets. An end-to-end model GndNet [136] integrates PointNet with a pillar feature encoding network for effective ground segmentation. Additionally, some approaches aim to predict traversability directly through CNNs; for example, Seo et al. [137] introduce a scalable framework for learning traversability from 3D LiDAR point clouds in a self-supervised fashion. Despite the promising results these learning-based methods have shown on various datasets, they often require substantial annotated datasets for training and struggle to adapt to different environments or LiDAR configurations, which poses limitations on their practical use.
@@
TABLE V: LiDAR-based traversability estimation methods for unstructured environments.
Type Reference Feature
Feature-based Thrun et al. [16] Min-Max Elevation
Lalonde et al. [127] EM
Ahtiainen et al. [128] SVM
Suger et al. [129] Random Forest Classifier
Lee and Chung [130] MLP
Guan et al. [131] Geometric Features
Paigwar et al. [136] PointNet
Seo et al. [137] Self-supervised Learning
Model-based Chen et al. [17] 2D GPR
Guizilini and Ramos [138] VHR
Shan et al. [139] BGK
Zhang et al. [140] MRF
Rummelhard et al. [141] CRF
Huang et al. [142] MRF
Forkel et al. [143] Recursive Gaussian State
Rodrigues et al. [144] B-spline
Cai et al. [145] Probabilistic
Fan et al. [146] 2.5D Traversability
Xue et al. [123] BGK
Xue et al. [147] Self-supervised Learning
@@
To address this challenge, Chen et al. [17] propose a method that transforms 2D GPR into a series of one-dimensional regressions, thereby reducing computational complexity. Nonetheless, this simplification tends to compromise the accuracy of the generated terrain model. An alternative approach, introduced in [138], utilizes Variational Hilbert Regression (VHR) for traversability analysis by projecting the point cloud into a Reproducing Kernel Hilbert Space (RKHS) for an approximate estimation of the terrain.
@@
In addition to Bayesian non-parametric inference, terrain modeling can also be approached using Markov Random Fields (MRF) or Conditional Random Fields (CRF). For instance, Zhang et al. [140] combine terrain smoothness assumptions with height measurements in a multi-label MRF framework, utilizing loopy belief propagation to achieve robust terrain modeling and ground segmentation.
@@
In another approach, Rummelhard et al. [141] model the terrain as a spatio-temporal CRF, estimating terrain elevation using the EM algorithm. Huang et al. [142] leverage the graph cut method to address a coarse-to-fine MRF model for effective ground segmentation. Moreover, Forkel et al. [143] present a probabilistic terrain estimation method framed as a recursive Gaussian state estimation problem, which they further extend into a dynamic resolution model in [150]. B-spline surfaces are also utilized in traversability analysis due to their excellent capability for fitting complex geometrical shapes, as demonstrated by [144]. Recent studies have also explored the integration of terrain models with vehicle kinematic models for traversability analysis. Cai et al. [145] introduce a probabilistic representation of traversability, modeling it as a distribution that is conditioned on the dynamic behaviors of the robot and the characteristics of the terrain. In a similar vein, Fan et al. [146] address uncertainties in risk-aware planning, proposing a 2.5D traversability evaluation method that factors in localization errors, sensor noise, vehicle dynamics, terrain characteristics, and various sources of traversability risk.
@@
Xue et al. [123] present a LiDAR-based terrain modeling approach that fuses multi-frame data to produce accurate and stable terrain models. Using normal distributions transform mapping, it fuses consecutive frames, and applies BGK and bilateral filtering to enhance stability while preserving sharp edges. Traversability is determined through geometric connectivity analysis of the terrain regions. Based on [123], Xue et al. [147] propose a self-supervised learning framework for terrain traversability in unstructured environments using contrastive label disambiguation. Pseudo labels are automatically generated by projecting driving experiences onto real-time terrain models. A prototype-based contrastive learning method refines these pseudo labels iteratively, eliminating ambiguities and enabling platform-specific traversability learning without human annotations.
III-B2 Vision-based Methods

### v4

@@ 2.2.3 Place Recognition / Relocalization
[187] Scan Context Multi-scene LiDAR
[145] Spatial Binary Pattern Orchard LiDAR
[240] Binary and Triangle Combined Descriptor Multi-scene LiDAR
[155] Global and Local Descriptor Natural Scene LiDAR/Camera
[141] Diverse Descriptors for Evaluation Dense Forest IMU/LiDAR
[6] PointNetPGAP-based Descriptor Horticulture LiDAR
@@
Place Recognition (PR) technology enables vehicles to identify previously visited environments, playing a crucial role in autonomous driving. This technology can be utilized to reduce drift errors of SLAM systems, ensuring reliable localization and globally consistent map construction. In unstructured environments, the PR task is particularly challenging due to the dynamic nature and sparse, diverse features of the surroundings. Such conditions often render methods [86, 85, 249] designed for urban environments ineffective. Currently, there are some studies that focus on PR technology in unstructured environments, as shown in Table 4.
@@
[187] propose STV-SC, which enhances the scan context [86] algorithm by incorporating segmentation and temporal verification. This method includes a range image-based 3D point segmentation algorithm to extract key structured information and a three-stage process for loop detection, significantly improving recall rates compared to state-of-the-art algorithms. [145] introduce a novel LiDAR-based PR algorithm that utilizes the spatial binary pattern to encode 3D spatial information into an eight-bit binary pattern. The two-stage hierarchical re-identification process with attention score maps further enhances the method’s effectiveness in orchard and public datasets. Another innovative method is [240], which combines binary and triangle descriptors to achieve robust 3D place recognition. This method forms triangles from keypoints and uses their side lengths to create a global descriptor invariant to pose changes, significantly improving precision and adaptability, especially in scenarios with large viewpoint variations.
@@
Further advancements in this domain involve integrating deep learning techniques for improved robustness in challenging scenarios. [155] present a method that integrates LiDAR and image data through a self-supervised 2D-3D feature matching process, leveraging deep learning for LiDAR feature extraction and relative pose estimation between point clouds. This approach improves robustness and accuracy in multi-robot relocalization tasks in unstructured environments. [141] evaluate handcrafted and learning-based models in dense forest environments, integrating the Logg3dNet model into a robust 6-DoF pose estimation system.
@@
[6] introduce PointNetPGAP-SLC, a model combining a global average pooling aggregator and a pairwise feature interaction aggregator to enhance descriptor generation. The Segment-Level Consistency (SLC) model used during training improves the robustness of descriptors. Evaluations on the dataset demonstrate that PointNetPGAP improves PR performance in horticultural settings by addressing descriptor ambiguity and leveraging the SLC model for enhanced training.
@@
Moreover, the development of specific datasets and evaluation frameworks has been crucial in advancing PR techniques in natural environments. [65] address the lack of benchmarks for natural, unstructured environments by introducing the Wild-Places [89] and WildScenes [194] datasets. These datasets enable comprehensive evaluation and development of localization and mapping algorithms in challenging unstructured settings, bridging the gap between urban and natural environments.
2.3 Odometry-based Methods
@@
[121] present Gamma-SLAM, an online stereo visual SLAM algorithm that uses a Rao-Blackwellized particle filter to estimate joint posterior distributions over poses and maps. This method incorporates visual odometry for accurate proposal distributions and represents maps through a Cartesian grid, maintaining a posterior distribution over elevation variance in each cell to effectively capture terrain features in unstructured environments. [244] introduce LOAM, a real-time LiDAR odometry and mapping method that divides the problem into high-frequency odometry and low-frequency mapping algorithms. This approach ensures low drift and low computational complexity, making it suitable for real-time applications. Extending this concept, [245] combine visual and LiDAR odometry (V-LOAM) to enhance localization and mapping by integrating high-frequency visual odometry for initial motion estimation and low-frequency LiDAR odometry for refinement, achieving robust performance even under aggressive motion and poor visual conditions.
@@
Further advancements in this domain include methods that focus on optimizing computational efficiency and robustness in varying terrains. [171] introduce LeGO-LOAM, a lightweight and ground-optimized LiDAR odometry and mapping method that achieves real-time six degree-of-freedom pose estimation using low-power embedded systems. This method leverages the ground plane for segmentation and optimization, enhancing accuracy and computational efficiency. Based on [171, 213], [175] establish fixed and generative road surface models to constrain the vertical drift errors of autonomous vehicles in outdoor scenarios. Since ground point clouds can be continuously and stably obtained, this ground-constrained approach could also be suitable for feature-sparse off-road environments.
@@
[53] present GPGM-SLAM, which leverages Gaussian process gradient maps to improve place recognition robustness and facilitate loop closure detection in planetary analog environments, demonstrating superior performance compared to traditional visual SLAM methods. Additionally, [18] propose a lightweight and efficient LiDAR odometry method that handles dense point clouds without significant preprocessing, using an adaptive keyframing system and a custom iterative closest point solver to reduce computational overhead.
@@
[241] introduce a novel voxel mapping method designed for LiDAR odometry, incorporating uncertainties from LiDAR measurement noise and pose estimation errors, and demonstrating superior accuracy and efficiency in various environments. [198] present a LiDAR-based SLAM algorithm that leverages neighborhood information constraints to enhance feature point discrimination and prevent mis-registration in rugged terrains, showing significant improvements in localization and mapping accuracy. [189] introduce X-ICP, which incorporates localizability-aware constraints in LiDAR registration to ensure reliable pose updates in geometrically uninformative environments. [25] propose an intensity-enhanced LiDAR SLAM framework that leverages both geometric and intensity information from LiDAR data to improve localization accuracy in sparse feature environments.
@@
The strategy of multi-sensor fusion has gained significant attention in recent years as an important means to effectively enhance the reliability and robustness of SLAM technology. For instance, [230] integrate IMU, wheel encoder, and LiDAR odometry to estimate vehicle motion, using IMU data to correct LiDAR scan distortions and provide better initial guesses for LiDAR odometry. This method demonstrates improved accuracy and high-frequency localization in various environments without prior information. Similarly, [246] present a modularized data processing pipeline that integrates range, vision, and inertial sensing to estimate motion and construct maps through multilayer optimization, ensuring high accuracy and low drift even under aggressive motion and environmental changes.
@@
[147] present LOCUS, a multi-sensor LiDAR-centric solution that integrates multiple sensing modalities and employs a robust scan-matching algorithm to ensure continuous operation despite sensor failures. This system demonstrates superior localization accuracy and robustness, particularly in challenging settings such as the DARPA Subterranean Challenge [33]. Additionally, [170] introduce LVI-SAM, a tightly-coupled LiDAR-visual-inertial odometry framework that leverages a factor graph for global optimization. This method enhances visual features with LiDAR depth information and employs a robust failure detection mechanism, achieving high accuracy and robustness in GPS-denied and feature-sparse environments.
@@
[97] present Switch-SLAM, which uses a switching structure to transition between LiDAR and visual odometry based on detected degeneracies, ensuring robust and accurate localization across various challenging environments without the need for heuristic tuning. Additionally, [92] introduce a multi-sensor fusion localization system that integrates LiDAR, vision, and IMU data to achieve high accuracy and robustness in GNSS-denied environments, dynamically adjusting fusion weights in response to sensor degradation and demonstrating superior performance in urban areas, tunnels, and unstructured mountainous roads.
Future research should investigate how mapping quality, including point cloud density and update frequency, affects pose estimation accuracy in unstructured terrains. For example, sparse LiDAR maps in low-feature environments such as deserts can lead to odometry drift and degraded localization, highlighting the need for adaptive strategies to mitigate map-induced errors.
3 Environmental Perception
@@
3.1 Challenges of Environmental Perception in Unstructured Environment
@@
3.2 LiDAR-based Methods
3.2.1 Traversability Estimation
@@
Traversability estimation methods based on geometric information focus on extracting and analyzing geometric features from LiDAR point cloud to classify the passability of the environment. These methods can be further categorized into feature-based methods and model-based methods, as shown in Table 5.
@@
Feature-based methods for analyzing traversability primarily rely on extracting characteristics from LiDAR point cloud to assess terrain passability. A notable example is the Min-Max elevation mapping technique introduced by [185]. This approach begins by projecting the 3D LiDAR point cloud into a 2D grid map. The method evaluates the traversability of each grid cell by examining the height variations of points within that cell. Besides height differences, other parameters such as average height [31], covariance [31], slope [123], and roughness[139] can also be employed to analyze traversability. These techniques are straightforward to implement and have relatively low computational demands. However, they often lack robustness and can struggle in adapting to environments characterized by rough terrain.
@@
To enhance robustness in traversability analysis, various machine learning algorithms can be employed. In [91], the Expectation Maximization (EM) algorithm is utilized to fit a Gaussian Mixture Model (GMM) based on labeled statistical features, categorizing terrains in forest environments into three distinct groups. A Support Vector Machine (SVM) classifier, as described in [3], distinguishes between traversable and non-traversable regions, with each grid cell represented in an NDT traversability map (NDT-TM). In unstructured environments, a random forest classifier has been developed in [180] to identify traversable areas. Furthermore, [96] detail the training of a Multi-Layer Perceptron (MLP) model for traversability mapping using a self-training algorithm. More recently, [55] combined geometric features, such as slope and step height, with semantic attributes to enhance traversability analysis.
@@
With the rise of deep learning, many researchers have focused on applying it for feature extraction and traversability analysis. To effectively leverage Convolutional Neural Networks (CNNs), two primary strategies have been proposed for processing unordered 3D point clouds. The first strategy involves re-encoding point clouds to create a dense input tensor suitable for CNNs. For instance, some studies have transformed point clouds into elevation maps [17], 2D Birds-Eye-View (BEV) maps [168], or range images [193], enabling CNNs to detect traversable regions in an end-to-end manner. The second strategy employs architectures like PointNet [152], designed specifically for unordered 3D point sets. An end-to-end model GndNet [146] integrates PointNet with a pillar feature encoding network for effective ground segmentation. Additionally, some approaches aim to predict traversability directly through CNNs; for example, [166] introduce a scalable framework for learning traversability from 3D LiDAR point clouds in a self-supervised fashion. Despite the promising results these learning-based methods have shown on various datasets, they often require substantial annotated datasets for training and struggle to adapt to different environments or LiDAR configurations, which poses limitations on their practical use.
@@
Table 5: LiDAR-based traversability estimation methods for unstructured environments.
Type Reference Description
Feature-based [185] Min-Max Elevation
[91] EM
[3] SVM
[180] Random Forest Classifier
[96] MLP
[55] Geometric Features
[146] PointNet
[166] Vehicle-terrain Interaction
Model-based [21] 2D GPR
[58] VHR
[172] BGK
[247] MRF
[161] CRF
[73] MRF
[40] Recursive Gaussian State
[158] B-spline
[14] Probabilistic
[37] 2.5D Traversability
[232] BGK
[233] Contrastive Label Disambiguation
@@
To address this challenge, [21] propose a method that transforms 2D GPR into a series of one-dimensional regressions, thereby reducing computational complexity. Nonetheless, this simplification tends to compromise the accuracy of the generated terrain model. An alternative approach, introduced in [58], utilizes Variational Hilbert Regression (VHR) for traversability analysis by projecting the point cloud into a Reproducing Kernel Hilbert Space (RKHS) for an approximate estimation of the terrain.
@@
In addition to Bayesian non-parametric inference, terrain modeling can also be approached using Markov Random Fields (MRF) or Conditional Random Fields (CRF). For instance, [247] combine terrain smoothness assumptions with height measurements in a multi-label MRF framework, utilizing loopy belief propagation to achieve robust terrain modeling and ground segmentation.
@@
In another approach, [161] model the terrain as a spatio-temporal CRF, estimating terrain elevation using the EM algorithm. [73] leverage the graph cut method to address a coarse-to-fine MRF model for effective ground segmentation. Moreover, [40] present a probabilistic terrain estimation method framed as a recursive Gaussian state estimation problem, which they further extend into a dynamic resolution model in [41]. B-spline surfaces are also utilized in traversability analysis due to their excellent capability for fitting complex geometrical shapes, as demonstrated by [158]. Recent studies have also explored the integration of terrain models with vehicle kinematic models for traversability analysis. [14] introduce a probabilistic representation of traversability, modeling it as a distribution that is conditioned on the dynamic behaviors of the robot and the characteristics of the terrain. In a similar vein, [37] address uncertainties in risk-aware planning, proposing a 2.5D traversability evaluation method that factors in localization errors, sensor noise, vehicle dynamics, terrain characteristics, and various sources of traversability risk.
@@
[232] present a LiDAR-based terrain modeling approach that fuses multi-frame data to produce accurate and stable terrain models. Using normal distributions transform mapping, it fuses consecutive frames, and applies BGK and bilateral filtering to enhance stability while preserving sharp edges. Traversability is determined through geometric connectivity analysis of the terrain regions. Based on [232], [233] propose a self-supervised learning framework for terrain traversability in unstructured environments using contrastive label disambiguation. Pseudo labels are automatically generated by projecting driving experiences onto real-time terrain models. A prototype-based contrastive learning method refines these pseudo labels iteratively, eliminating ambiguities and enabling platform-specific traversability learning without human annotations.
3.2.2 Semantic Segmentation